OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4

OpenStudy (anonymous):

Which of the following is a correct next step to solve the equation? x + 2 = -4 -x - 2 = 4 x + 2 = 4 x - 2 = -4

OpenStudy (anonymous):

I think the answer to this question is the last option. @phi @tkhunny @e.mccormick @ManImStumped @ohh_munoz @One098

OpenStudy (anonymous):

@Abhisar @goformit100 @Compassionate

OpenStudy (anonymous):

So the third one?

OpenStudy (anonymous):

thnx

OpenStudy (phi):

You were correct the first time. | x | = 4 means either x = 4 or -x = 4 , which means x= -4 if you used the 3rd option x + 2 = 4 and solve for x (add -2 to both sides) we get x+2 -2 = 4 - 2 x +0 = 2 x=2 now test in the original |x - 2| = 4 replace x with 2 and we get | 2-2 | = 4 ? |0| = 4 ? nope. that shows the 3rd option is not correct.

OpenStudy (jojokiw3):

Right. I'm sorry.

OpenStudy (anonymous):

Yeah, I found out. THANK YOU SO MUCH! and @jojokiw3 is ok. :)